# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Introduction
9:   $b(\bm{X})^{-}$ $\triangleright$ Smallest set (given metric $m$) that yields a solution to Problem (5) when $g(\bm{X})=b(\bm{X})^{-}$
10:   $\{w^{*}\}_{i=1}^{n}$ $\triangleright$ Resulting weights from (5) with $g(\bm{X})=b(\bm{X})^{-}$
From a design-based perspective, selecting the covariate functions that are each conditionally the most strongly predictive of treatment will tend to result in the smallest $g(\bm{x})$ set that yields a solution to Problem (5). While selecting this set may not minimize the design-based components of estimator error, this set does tend to result in smaller design-based estimator error and a simpler interpretation of the resulting estimator. Thus, for the design-based metric, we propose and implement the Spearman semipartial correlation; in step four the maximum correlation is used to identify $b^{*}(\bm{x})$. We choose this metric because it is rank-based and reflects the impact of removing a covariate in a hypothetical propensity score model, however other metrics could work well here. Our design-based estimator takes the same form as Equation (4) with weights computed with the $g(\bm{x})$ set selected by Algorithm 1.
From a model-based perspective, selecting the covariate functions that do not contribute to treatment effect heterogeneity as the $g(\bm{x})$ set will tend to minimize estimator error. Thus, we implement the absolute value of the one-step doubly robust estimator of treatment effect modification, proposed by Boileau et al., (2025), as the model-based metric to use within Algorithm 1; in step four the minimum of this metric is used to identify $b^{*}(\bm{X})$. Our model-based estimator is a cross-fit style estimator to preserve downstream inference. We provide further discussion about this metric and model-based estimator in supplementary Section 4.2. For either metric/estimator, Algorithm 1 can be initiated with a non-empty $b(\bm{x})^{-}$ set (step 2) if there are covariates that one wants to include in the $g(\bm{x})$ set a priori. If minimizing error with respect to the ATE is not a priority, $g(\bm{x})$ could be chosen such that the modified $g(\bm{x})$ population is of scientific interest which we discuss further in supplementary Section 4.1.
@@
We generate a variety of simulation scenarios that vary in the number of confounders (i.e., covariates), level of overlap, level of treatment effect heterogeneity, and proportion of covariates that are effect modifiers. Through these scenarios, we first aim to validate the estimator corresponding to the weights derived through Problem (5) when the true $\tau(\bm{x})$ function is known. In addition, we seek to determine in which data generating scenarios the proposed methods for a completely data-driven selection of $g(\bm{x})$ in Section 6 yield estimators with improved performance when compared to IPW ATE estimation, minimal weights, and the ATO when there is no solution to Problem (3).
@@
For clarity in presentation, we drop the observation index on the variables. For $p=20$ and $p=100$ we simulate $V_{1}$ - $V_{p}$ from a truncated normal distribution where $E[V_{i}]=0$, $\text{Var}(V_{i})=1$ and $\text{Cov}(V_{i},V_{j})=0$ for all $i=1,\ldots,p$ and $i\neq j$. To obtain a mix of continuous and categorial coefficients, we take $X_{1}$ - $X_{p/2}=$ $V_{1}$ - $V_{p/2}$ and $X_{j}=I[V_{j}<0]$ for $j=p/2+1,\ldots p$. We generate the true propensity scores with a logistic model, $e(\bm{X})=\{1-\exp(-\sum_{i=0}^{p}\alpha_{i})\}^{-1}$. In this model, $\alpha_{0}$ is selected to obtain the desired percent of treated observations (20%, 40%) and all other coefficients are determined as described in supplementary Section 5. We simulate treatment assignment as $Z\sim\text{Bernoulli}\{e(\bm{X})\}$ and the outcome as $Y=\mu_{z}(\bm{X})+N(0,1)$ for $Z=z$. Both $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ are modeled as linear combinations of $X_{1},\ldots,X_{p}$. For $\theta=25\%,75\%$ of covariates, their corresponding $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ coefficients are the equal such that these covariates are not effect modifiers. For all other covariates, the $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ coefficients are determined as described in supplementary Section 5
We generate a total of 72 simulation scenarios that are all combinations $p=20,100$, 20% and 40% treated, 25% and 75% of covariates being effect modifiers, three levels of overlap, and three levels of treatment effect heterogeneity. For each simulation scenario we generate 1,000 independent datasets of size either 1,000 ($p=20$) or 2,000 ($p=100$) observations. We calculate the true value of the ATE using 10 million Monte Carlo samples. We compute 1) IPW, direct balancing weights, and minimal weights estimators for the ATE; 2) our proposed estimator given the true $g(\bm{X})$ and the design- and model-based estimators proposed in Section 6; and 3) the IPW and balancing weights (i.e., all covariates in the set $g(\bm{X})$) estimators for the ATO. All balancing weights are computed with the stable balancing weights deviance measure (Zubizarreta,, 2015). We compare the mean squared error (MSE) of these estimators with respect to the true ATE.
@@ Results
Our proposed estimator implemented with the true $\tau(\bm{x})$ function has the minimum estimator MSE with respect to the ATE across almost all simulation scenarios where there is a solution to Problem (5) for the true $\tau(\bm{x})$ function (Figure 2 and supplementary Figure 4). Specifically, this estimator tends to be minimally biased and have comparable variance to the minimal weights estimator and ATO estimator (supplementary Figures 5-8). Thus, when there is a lack of overlap and no solution to Problem (3), our proposed estimator is the only estimator that has minimal bias for the ATE; both the minimal weights and the ATO estimators tend to be substantially biased for the ATE. However, it is rare that the true $\tau(\bm{x})$ is known, so for the remainder of this section we focus on performance of our proposed design- and model-based estimators.
@@
Figure 2: Estimator mean squared error (MSE) for all 20% treated simulation scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing optimization problem did not exist for all 1,000 datasets. The horizontal axis indicates levels of treatment effect heterogeneity while each panel is labeled by different overlap levels, number of predictors, and percent of covariates that are effect modifiers.
@@
Across almost all simulation scenarios, either our proposed design- or model-based estimator has minimum estimator MSE with respect to the ATE compared to minimal weights or either ATO estimator (Figure 2 and supplementary Figure 4). Furthermore, the design-based estimator has lower MSE than the IPW ATE estimator across all scenarios, which none of the other estimators achieve. The minimal weights estimator tends to have the largest MSE, even larger than ATE and ATO estimators, when there is moderate-high treatment effect heterogeneity and low overlap. The model-based estimator tends to perform better than the design-based estimator for $p=20$, in terms of MSE, while the reverse is true when $p=100$. We likely see this trend because there may not be sufficient sample size for estimating treatment effect modification and computing the weights with sample splitting. Thus, the model-based estimator requires a substantially larger sample size for high dimensional covariates for comparable performance to the design-based estimator. We see similar trends in estimator MSE between 20% treated (Figure 2) and 40% treated scenarios (supplementary Figure 4), though for 40% treated scenarios all estimators substantially outperform the IPW estimator in terms of MSE.
The scenarios where the minimal weights estimator tends to have lower MSE than either of our proposed estimators are ones with lower treatment effect heterogeneity (in terms of both the scale of covariate-level treatment effect heterogeneity and the % of covariates that are effect modifiers). In these scenarios, the design- and model-based estimators tend to have larger estimator variance, but smaller statistical bias with respect to the ATE compared to the minimal weights estimator (supplementary Figure 5). While the small variance of the minimal weights estimator is desirable, the estimator is substantially biased for the ATE such that the corresponding 95% CI may not include the true ATE. Furthermore, there is no clear causal estimand for which the minimal weights estimator is consistent. Thus, even in scenarios in which the minimal weights estimator has lower MSE, our proposed estimators may be preferable.
@@
We apply our weighting method for ATE estimation to observational data from the MIMIC-III v1.4 critical care database (Johnson et al.,, 2016). Specifically, we perform a reanalysis of Hsu et al., (2015) which examines effect of indwelling arterial catheters (IACs) on mortality in patients with respiratory failure. The data can be obtained with queries provided at the GitHub repository https://github.com/MIT-LCP/mimic-code. The data has 2,522 observations of mechanically ventilated patients, where 51.5% of patients received IAC. The outcome of interest is an indicator of mortality 28 days within hospital admission. Pre-treatment covariates that may be confounders include demographic information, baseline measurements (e.g., blood pressure, lab values), risk scores, and more. Missing data is imputed with single imputation and we include missing data indicators in the set of pre-treatment covariates. In total, we identify 72 pre-treatment covariates of interest.
@@
There is a lack of overlap in the estimated propensity scores in this data and there are many propensity scores close to 1 (supplementary Figure 9). Due to this lack of overlap, there is no solution to Problem (3) and it is difficult to achieve sufficient covariate balance. We estimate the ATE with IPW, minimal weights, and our proposed design- and model-based estimators that use Algorithm (1) to select $g(\bm{x})$. Since rare binary covariates tend to be challenging to exactly balance, we also compute our proposed estimators when initializing Algorithm (1) with the set of rare binary covariates with outcomes that occur for < 5% of observations. We also estimate the ATO using IPW and balancing weights (i.e., all covariates in the set $g(\bm{x})$). We then compare the covariate balance, treatment effect estimates, and Wald-type bootstrapped 95% confidence intervals (CIs) between all methods.
@@
The treatment effect estimates and 95% confidence CIs are similar for all the estimators except the IPW estimator, which corresponds to a more negative estimate and a larger confidence interval (Figure 3). However, all estimates and 95% CIs indicate that IAC non-significantly reduces 28 day mortality by a small percentage. The fact that the minimal weights, ATO, and our proposed estimates are similar likely indicates that there is limited treatment effect heterogeneity. Thus, the main difference between these estimators for this application are in terms of covariate balance and target population. Figure 3 shows the weighted treated and control sample for three pre-treatment covariates in comparison to the sample mean for those covariates (the black horizontal line). IPW and minimal weights do not exactly balance any of the covariates, with corresponding average standardized mean differences (SMD) of $0.161$ and $0.008$, respectively. In contrast, all other weights exactly balance the treated and control distributions. Our proposed design-based weights where Algorithm 1 is initiated with rare binary covariates corresponds to the smallest $g(\bm{x})$ set. In Figure 3, these weights balance both diastolic pressure and the SOFA score to the target population, which is not true of any of the other estimators. Supplementary Figures 10-12 show the SMD for all covariates 1) between treated and control groups; and 2) between treated/control groups and the target population. While these estimators yield similar estimates and 95% confidence intervals for this application, our proposed estimators tend to have improved covariate balance and/or correspond to a minimally modified target population.
@@
Figure 3: ATE estimates and 95% confidence intervals of the effect of IAC on 28-day mortality (bottom figure) and weighted covariate balance of three covariates (three top figures) for eight different weighting/estimation methods (indicated by color). For the top three figures, the horizontal line indicates the sample mean of the covariates in the target population and vertical lines indicate the extent of covariate imbalance for each method.
@@
Transporting “health care hotspotting” RCT to a Midwestern U.S. academic health center population
A recent intervention of popular interest, termed ‘health care hotspotting’, aimed to improve healthcare delivery for individuals who interact substantially with the healthcare system. Despite the promising nature of this intervention, a randomized controlled trial (RCT) in Camden, NJ found a null effect of this intervention on readmission rates Finkelstein et al., (2020). However, a re-anlaysis of this trial by Yang et al., (2023) found a significant effect of this intervention in the population of individuals who had a high probability of engagement with the randomized intervention. Since there is evidence that the treatment effect varies by population, we are interested in determining the effect of this intervention in a Midwestern U.S. academic health center population.
@@
In the RCT, individuals were randomized to standard of care or an intensive, targeted follow-up after admission. The outcome of interest is a binary indicator of readmission within 30 days. We include all pre-treatment covariates that are in common between both datasets, which includes demographics and hospital stay characteristics, for a total of 21 covariates. We perform a complete case analysis (< 30 missing observations) with a total of 781 trial population observations and 1305 target population observations. There is minimal overlap in the estimated probability of being in the target population (supplementary Figure 13). However, the primary driver of this lack of overlap is race; the trial populations is 84.9% non-white while the target population is only 13.6% non-white. When race is removed from the model predicting the probability of being in the target population, there is substantial overlap (supplementary Figure 13). Thus, we implement our proposed design-based estimator and our weighted estimator when $g(\bm{x})$ a priori only includes race. We compare this estimator to the normalized inverse odds weighted (IOW) estimator (Dahabreh et al.,, 2020), the inverse odds weighted estimator where population probabilities less than $0.1$ and more than $0.9$ are trimmed (Crump et al.,, 2009), and the standard stable balancing weights estimator. We report covariate balance, transported treatment effect estimates, and and Wald-type bootstrapped 95% CIs for these methods.
@@
The trimmed estimator has the largest treatment effect estimate, but also the largest 95% CI, likely due to the small sample size after trimming (Figure 4). The IOW and balancing weights estimates and 95% CIs indicate a non-significant reduction in 30-day readmission; however, there was no solution to the direct balancing weights optimization problem for many of the bootstrapped datasets such that a true 95% bootstrapped CI cannot be derived. While the 95% CIs for both of our proposed estimators do not include zero, the interval is substantially smaller and the estimates indicate that the intervention may actually increase hospital readmission rates in the target population. Supplementary Figures 14-16 show the SMD for all covariates 1) between treated and control groups; and 2) between treated/control groups and the target population. Both the IOW and trimmed estimators achieve poor covariate balance with average SMDs of $0.140$ and $0.382$. Balancing weights and our proposed weights achieve exact covariate balance; however, our proposed weights do modify the sample mean of the race covariate such that the estimate applies to a population that is approximately 80% non-white. Therefore, our proposed estimator indicates that the intervention may in fact be detrimental for 30-day readmission rates for this slightly modified target population and for original target population when race is not a treatment effect modifier.
@@
Figure 4: Transported treatment effect estimates and 95% confidence intervals of the effect of ‘health care hotspotting’ on 30-day readmission rates for a Midwestern healthcare center (bottom figure) and weighted covariate balance of two covariates (two top figures) for for different weighting/estimation methods (indicated by color). For the top two figures, the horizontal line indicates the sample mean of the covariates in the target population and vertical lines indicate the extent of covariate imbalance for each method.

### v2

@@ Introduction
9:   $b(\bm{X})^{-}$ $\triangleright$ Smallest set that yields a solution to Problem (5) when $g(\bm{X})=b(\bm{X})^{-}$
10:   $\{w^{*}_{i}\}_{i=1}^{n}$ $\triangleright$ Resulting weights from Problem (5) with $g(\bm{X})=b(\bm{X})^{-}$
From a design-based perspective, selecting the covariate functions that are each conditionally the most predictive of treatment will tend to result in the smallest $g(\bm{x})$ set that yields a solution to Problem (5). While selecting this set may not minimize the design-based simplification of estimator error, this set does tend to result in smaller design-based estimator error and simpler estimator interpretation. Thus, for the design-based metric, we propose and implement Spearman semipartial correlation; in step four, the maximum correlation is used to identify $b^{*}(\bm{x})$. We choose this metric because it is rank-based and reflects the impact of removing a covariate in a hypothetical propensity score model, however other metrics could work well here. Our design-based estimator then takes the same form as Equation (4) with weights computed with the $g(\bm{x})$ set selected by Algorithm 1.
From a model-based perspective, selecting the covariate functions that do not modify the treatment effect as $g(\bm{x})$ will tend to minimize estimator error. Thus, we implement the absolute value of the one-step doubly robust estimator of treatment effect modification, proposed by Boileau et al. (2025), as the model-based metric to use within Algorithm 1; in step four, the minimum of this metric is used to identify $b^{*}(\bm{x})$. Our model-based estimator is a cross-fit style estimator to preserve downstream inference; we provide further discussion of this model-based metric and estimator in supplementary Section 5.2. For either metric/estimator, Algorithm 1 can be initiated with a non-empty $b(\bm{x})^{-}$ set (step 2) if there are covariates that one wants to include in the $g(\bm{x})$ set a priori. If minimizing error with respect to the ATE is not a priority, $g(\bm{x})$ could be chosen such that the modified $g(\bm{x})$ population is of scientific interest which we discuss further in supplementary Section 5.1.
@@
We generate a variety of simulation scenarios that vary in the number of confounders (i.e., covariates), level of overlap, level of treatment effect heterogeneity, and proportion of covariates that are effect modifiers. Through these scenarios, we first aim to validate the estimator corresponding to the weights derived through Problem (5) when the true $\tau(\bm{x})$ function is known. In addition, we seek to determine in which data generating scenarios the proposed methods for a completely data-driven selection of $g(\bm{x})$ in Section 5 yield estimators with improved performance when compared to IPW, minimal weights, and overlap weights estimators when there is no solution to Problem (3).
@@
For clarity in presentation, we drop the observation index on the variables. For $p=20$ and $p=100$ we simulate $V_{1}$ - $V_{p}$ from a truncated normal distribution where $E[V_{i}]=0$, $\text{Var}(V_{i})=1$ and $\text{Cov}(V_{i},V_{j})=0$ for all $i=1,\ldots,p$ and $i\neq j$. To obtain a mix of continuous and categorial coefficients, we take $X_{1}$ - $X_{p/2}=$ $V_{1}$ - $V_{p/2}$ and $X_{j}=I[V_{j}<0]$ for $j=p/2+1,\ldots p$. We generate the true propensity scores with a logistic model, $e(\bm{X})=\{1-\exp(-\sum_{i=0}^{p}\alpha_{i})\}^{-1}$. In this model, $\alpha_{0}$ is selected to obtain the desired percent of treated observations (20%, 40%) and all other coefficients are determined as described in supplementary Section 6. We simulate treatment assignment as $Z\sim\text{Bernoulli}\{e(\bm{X})\}$ and the outcome as $Y=\mu_{z}(\bm{X})+N(0,1)$ for $Z=z$. Both $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ are modeled as linear combinations of $X_{1},\ldots,X_{p}$. For $\theta=25\%,75\%$ of covariates, the corresponding $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ coefficients are equal (i.e., these covariates are not effect modifiers). The remaining $\mu_{0}(\bm{x})$ and $\mu_{1}(\bm{x})$ coefficients are determined as described in supplementary Section 6.
We generate a total of 72 simulation scenarios of all combinations $p=20,100$, 20% and 40% treated, 25% and 75% of covariates being effect modifiers, three levels of overlap, and three levels of treatment effect heterogeneity. For each simulation scenario we generate 1,000 independent datasets with either $n=1000$ ($p=20$) or $n=2000$ ($p=100$) observations. We compute 1) IPW, direct balancing weights, and minimal weights estimators for the ATE; 2) our proposed estimator given the true $g(\bm{X})$ and the design- and model-based estimators proposed in Section 5; and 3) the IPW and balancing weights (i.e., all covariates in the set $g(\bm{x})$) estimators for the ATO. All balancing weights are computed with the stable balancing weights deviance measure (Zubizarreta, 2015) and with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$. We compare the mean squared error (MSE) of these estimators with respect to the true ATE.
@@ Results
Our proposed estimator implemented with the true $\tau(\bm{x})$ function has the minimum estimator MSE with respect to the ATE across almost all simulation scenarios where there is a solution to Problem (5) for the true $\tau(\bm{x})$ function (Figure 2 and supplementary Figure 7). Specifically, this estimator tends to be minimally biased and have comparable variance to the minimal weights and ATO estimators (supplementary Figures 8-11). Furthermore, when there is a lack of overlap and no solution to Problem (3), our proposed estimator is the only estimator that has minimal bias for the ATE. However, it is rare that the true $\tau(\bm{x})$ is known, so we focus on the performance of our proposed design- and model-based estimators that use a data-driven procedure to select $g(\bm{x}).$
@@
Figure 2: Estimator mean squared error (MSE) for all 20% treated simulation scenarios. The line indicates the MSE for the IPW ATE estimator for each scenario. The asterisk, *, indicates estimators where a solution to the balancing optimization problem did not exist for all datasets. The horizontal axis indicates levels of treatment effect heterogeneity while each panel is labeled by different overlap levels, number of predictors, and percent of covariates that are treatment effect modifiers.
@@
Across almost all simulation scenarios, either our proposed design- or model-based estimator has minimum estimator MSE with respect to the ATE compared to either the minimal weights or ATO estimators (Figure 2 and supplementary Figure 7). Furthermore, the design-based estimator is the only estimator that has 1) a solution to Problem (5); and 2) lower MSE than the IPW ATE estimator across all scenarios. The minimal weights estimator tends to have the largest MSE, even larger than the ATE and ATO estimators, when there is moderate-high treatment effect heterogeneity and low overlap. When there is low treatment effect heterogeneity, the minimal weights estimator tends to have the smallest MSE. However, this is due to the estimator achieving minimal variance but possessing substantial statistical bias; thus, our proposed estimators may still be preferable in these scenarios (supplementary Figure 8). Our model-based estimator tends to perform better than the design-based estimator for $p=20$, while the reverse is true when $p=100$. We likely see this trend because there may not be sufficient sample size for estimating treatment effect modification and computing the weights with sample splitting for $p=100$. We see similar trends in estimator MSE between the 20% treated (Figure 2) and 40% treated scenarios (supplementary Figure 7), though for the 40% treated scenarios all estimators substantially outperform the IPW estimator.
Two illustrative applications
We apply our proposed weighted estimators for ATE estimation to an EHR study and adapt Problem (5) for transportability to transport a treatment effect from an RCT to a Midwestern academic health center population. We compare our proposed estimators to IPW, direct balancing, and minimal weights estimators for the ATE and ATO and demonstrate their superior performance in terms of covariate balance and estimator variance for these two applications.
@@
We perform a reanalysis of Hsu et al. (2015) which estimates the ATE of indwelling arterial catheters (IACs) on mortality in patients with respiratory failure using observational data from the MIMIC-III v1.4 critical care database (Johnson et al., 2016). The data has 2,522 observations of mechanically ventilated patients, where 51.5% of patients received IAC. The outcome of interest is an indicator of mortality 28 days within hospital admission. Missing data is imputed with single imputation. Pre-treatment covariates include demographic information, baseline measurements (e.g., blood pressure, lab values), risk scores, and missing data indicators. In total, we identify 72 pre-treatment covariates of interest. There is a lack of overlap in the estimated propensity scores in this data with many propensity scores close to 1 (supplementary Figure 12). Due to this lack of overlap, there is no solution to Problem (3) with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$. We estimate the ATE with IPW, minimal weights, and our proposed design- and model-based estimators that use Algorithm (1) to select $g(\bm{x})$. Since rare binary covariates tend to be challenging to exactly balance, we also compute our proposed estimators when initializing Algorithm (1) with the set of binary covariates whose outcomes occur for < 5% of observations. We also estimate the ATO using IPW and balancing weights (i.e., all covariates are in the $g(\bm{x})$ set). All balancing weights are computed with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$. We compare the covariate balance, treatment effect estimates, and Wald-type bootstrapped 95% confidence intervals (CIs) between all methods.
@@
The treatment effect estimates and 95% CIs are similar for all estimators except the IPW estimator, which corresponds to a more negative estimate and a larger 95% CI (Figure 3). However, all estimates and 95% CIs indicate that IAC non-significantly reduces 28 day mortality by a small percentage. The fact that the minimal weights, overlap weights, and PRTBW estimates are similar likely indicates that there is limited treatment effect heterogeneity. Thus, these estimators primarily differ in their target populations and the covariate balance achieved. Figure 3 shows the weighted treated and control sample means for three covariates in comparison to the sample mean (the black horizontal line). IPW and minimal weights do not exactly balance any of the covariates, with corresponding average standardized mean differences (SMD) of $0.161$ and $0.008$, respectively. In contrast, all other weights exactly balance the treated and control distributions. The smallest $g(\bm{x})$ set is selected for our method when Algorithm 1 is initiated with rare binary covariates. In Figure 3, the corresponding design-based weights balance both diastolic pressure and the SOFA score to the target population, which is not true of any of the other weights. Supplementary Figures 13-15 show the SMD for all covariates 1) between treated and control groups; and 2) between treated/control groups and the target population. While these estimators yield similar estimates and 95% CIs for this application, our proposed estimators tend to have improved covariate balance and correspond to a minimally modified target population.
@@
Figure 3: ATE estimates and 95% confidence intervals of the effect of IAC on 28-day mortality (bottom figure) and weighted covariate balance of three covariates (top figures) for eight different weighting/estimation methods. For the top figures, the horizontal line indicates the sample mean of the covariates in the target population and vertical lines indicate the extent of covariate imbalance for each method.
@@
Transporting a “health care hotspotting” RCT to a Midwestern U.S. academic health center population
@@
A recent intervention of popular interest, termed “health care hotspotting”, aimed to improve healthcare delivery for individuals who interact substantially with the healthcare system through intensive, targeted follow-up after admission. Despite the promising nature of this intervention, an RCT in Camden, NJ found a null effect of this intervention on readmission rates (Finkelstein et al., 2020). However, there is evidence that the treatment effect varies by population (Yang et al., 2023) such that we are interested in determining the effect of this intervention in a Midwestern U.S. academic health center population. We perform a complete case analysis (< 30 missing observations) with a total of 781 trial population observations and 1305 target population observations. The outcome of interest is a readmission within 30 days. We include all common pre-treatment covariates between both datasets, which includes demographics and hospital stay characteristics, for a total of 21 covariates. There is minimal overlap in the estimated probability of being in the target population primarily due to race; the trial populations is 84.9% non-white while the target population is 13.6% non-white (supplementary Figure 16). Thus, we implement our proposed design-based estimator and our weighted estimator when $g(\bm{x})$ a priori only includes race. We compare this estimator to the normalized inverse odds weighted (IOW) estimator (Dahabreh et al., 2020), the IOW estimator with population probabilities outside of $[0.1,0.9]$ trimmed (Crump et al., 2009), and the direct balancing weights estimator. All balancing weights are computed with the additional constraints of $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$. We report covariate balance, transported treatment effect estimates, and Wald-type bootstrapped 95% CIs for these methods.
@@
The IOW and direct balancing weights estimates and 95% CIs indicate a non-significant reduction in 30-day readmission (Figure 4); however, there is no solution to the direct balancing weights optimization problem for many of the bootstrapped datasets such that a true 95% bootstrapped CI cannot be derived. The trimmed estimator has the largest treatment effect estimate, but also the largest 95% CI, likely due to the small sample size after trimming. While the 95% CIs for both of our proposed estimators include zero, the intervals are substantially smaller and the estimates indicate that the intervention may actually increase hospital readmission rates in the target population. Both the IOW and trimmed estimators achieve poor covariate balance with average SMDs of $0.140$ and $0.382$. Supplementary Figures 17-19 show the SMD for all covariates 1) between treated and control groups; and 2) between treated/control groups and the target population. Direct balancing weights and our proposed weights achieve exact covariate balance; however, our proposed weights do modify the sample mean of the race covariate such that the estimate applies to a population that is approximately 80% non-white. Therefore, the PRTBW estimates indicate that the intervention may in fact increase 30-day readmission rates for either 1) this slightly modified target population; or 2) for the original target population when race is not a treatment effect modifier.
@@
Figure 4: Transported treatment effect estimates and 95% confidence intervals of the effect of “health care hotspotting” on 30-day readmission rates for a Midwestern healthcare center population (bottom figure) and weighted covariate balance of two covariates (top figures) for five different weighting/estimation methods. For the top figures, the horizontal line indicates the sample mean of the covariates in the target population and vertical lines indicate the extent of covariate imbalance for each method.